OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in simplifying the quantity of two x squared plus eight x minus forty-two, divided by the quantity of x squared plus twelve x plus thirty five Be sure to list restrictions.

OpenStudy (anonymous):

\[\frac{ 2x ^{2}+8x-42 }{ x ^{2}+12x+35 }\] factor each of them \[\frac{ (2x-6)(x+7) }{ (x+7)(x+5) }\] thus = \[\frac{ 2x-6 }{ x+5 }\]